OpenStudy (anonymous):

what is the y intercept 3(y-5)=6x+8

TheSmartOne (thesmartone):

y intercept means x=0. So plug in x=0 and find y.

TheSmartOne (thesmartone):

So 3(y-5)=6(0)+8 And solve for y.

TheSmartOne (thesmartone):

Have you gotten anything?

OpenStudy (anonymous):

not yet

TheSmartOne (thesmartone):

Remember anything times 0 equals to 0

TheSmartOne (thesmartone):

Anything now?

TheSmartOne (thesmartone):

Also read this tutorial for new OS users. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

OpenStudy (anonymous):

3y-15=8 ? I'm lost

TheSmartOne (thesmartone):

Correct

TheSmartOne (thesmartone):

3(y-5)=3y-15 6*0+8=0+8=8 So 3y-15=8 Now just add 15 on both sides and then divide by 3.